# Multiple References with Meaningful Variations Improve Literary Machine Translation

### v1

@@ 6.4 Performance difference between languages
All source languages in our datasets are also in the mT5-large and LLaMA-2-7B training data. However, as described in Section 5.1, non-English languages are only of an extremely small percentage of the total training data.
We compare the gain between Unfiltered All and Medium+High All among different languages under different metrics in Figure 3, and we also compare this gain with respect to the amount of fine-tuning data in Unfiltered All.
While among different metrics, Medium+High All outperforms Unfiltered All (Table 2), under the same model, each language’s metric gain is different. Surprisingly, for both mT5-large and LLaMA-2-7B, more fine-tuning data for a language (e.g. French and Russian have more fine-tuning data) does not result in proportionally more gain.
The gain for languages with less fine-tuning data also seems inconsistent between mT5-large and LLaMA-2-7B in both the amount and direction, e.g. Classical Chinese mostly has negative gain for mT5-large while much bigger positive gain using LLaMA-2-7B. However, overall, between metrics, the direction of gain for a language under the same model, is consistent.
@@ 6.5 Ablation study on the highly similar paraphrases
The Medium+High datasets (Medium+High 10%, 20% and ALL) are only different by the quantity of high semantic similarity paraphrases (Table 2 and Section 6.2,), in this ablation study, we further examine the characteristics of these high semantic similarity paraphrases and their impact on MT performance. We classify these high semantic similarity paraphrases into 3 categories:
@@
For each category above, we combine Medium All and 10% additional paraphrases from that category and create a new dataset. These new datasets have the same number of source texts, #source = 57,987), with the exception that the first category (sim_p = 1.0) does not have enough data. To counter that, since the paraphrases in the first category are pure repetition, we create more training instances by randomly selecting source texts from the other two categories and repeating one of its reference translations.
Across different metrics, mT5 fine-tuned on Medium+1.0 provides the best performance, while LLaMA-2-7B performs best with Medium+0.85-0.9 (Table 3). We tested the statistical significance (p-value < 0.05) of the results in Table 3 using the paired bootstrap resampling method (Koehn, 2004).
Given the conflicting results from the two LLMs in this ablation study, it is inconclusive what the best kind of high semantic similarity is for achieving the best MT performance. We recommend simply including all medium and high semantic similarity paraphrases (Medium+High All) as suggested in Section 6.2.
@@ 7 Conclusion
Unlike more utilitarian domains of translation, literary works encourages multiple translators to try their hand at improving on their predecessors. To make the most of this valuable resource for training MT systems, we explore the optimal way to employ multiple references when fine-tuning LLMs for machine translation.
@@
Additionally, we observe that when holding the total number of training instances constant, single-reference (having more source texts with one target text each) is slightly better than multiple-reference (less source text with multiple target texts each). Future research could test these findings in other MT domains.
@@ 9 Potential Risks
Appendix A Medium All dataset statistics
@@
Portuguese 4 1021 3261 0.313
Swedish 3 1249 2591 0.482
Italian 2 1325 1912 0.693
Czech 4 1759 2886 0.609
Japanese 9 1779 2483 0.716
Chinese 7 2199 3671 0.599
Norwegian 2 3011 4756 0.633
German 16 6136 8905 0.689
Russian 27 20003 34030 0.588
French 32 32353 51052 0.634
Table 4: Composition of the Medium All dataset. "#src" is the total number of unique source paragraphs selected, while column "total" is the all unfiltered number of paragraphs. "%used" is "#src" divided by "total".
@@ Appendix B Example of an untranslatable word
Appendix C Example source texts and target texts from Par3 dataset
source language source text human references
@@ Appendix C Example source texts and target texts from Par3 dataset
Table 5: Examples of source texts and human references from the Medium All dataset where there are meaningful variations between human references for each source text.

### v2

@@ 6.4 Performance difference between languages
We compare the performance gain between the Unfiltered and Medium+High (#source = 68k) datasets across different languages and metrics, as shown in Figure 3 in the Appendix. Additionally, we evaluate this gain in relation to the amount of fine-tuning data in the Unfiltered dataset.
While among different metrics, Medium+High outperforms Unfiltered (Table 3), under the same model, each language’s metric gain is different. Surprisingly, for all models, more fine-tuning data for a language (e.g., French and Russian have more fine-tuning data) does not result in proportionally more gain.
The gain for languages with less fine-tuning data also seems inconsistent between different models in both the amount and direction, e.g., Classical Chinese mostly has negative gain for mT5-large while much bigger positive gain using LLaMA-2-7B and Opus-MT. However, overall, between metrics, the direction of gain for a language under the same model, is consistent.
@@ 6.5 Ablation study on the highly similar paraphrases
The Medium+High datasets in Table 3 are only different by the quantity of high semantic similarity paraphrases. In this ablation study, we further examine the characteristics of these high semantic similarity paraphrases and their impact on MT performance. We classify these high semantic similarity paraphrases into three categories:
@@
For each category, we combine the full Medium dataset (#source = 50k) with 10% additional paraphrases from the same category to form a new dataset (#source = 58k). However, the first category (sim_p = 1.0) does not have enough additional instances. To counter that, since the paraphrases in the first category are pure repetition, we create more training instances by randomly selecting source texts from the other two categories and repeating one of its reference translations.
Across different metrics, Medium+0.85–0.9 demonstrates a marginally improved performance with the caveat that under mT5, performance across these different datasets is similar. We tested the statistical significance (p-value < 0.01) of the results in Table 4 with Medium+1.0 as the baseline.
@@ 7 Conclusion
Unlike more utilitarian domains of translation, literary works encourages multiple translators to try their hand at improving on their predecessors. To make the most of this valuable resource for training MT systems, we explore the optimal way to employ multiple references when fine-tuning different models for machine translation.
@@
Additionally, we observe that when holding the total number of training instances constant, single-reference (having more source texts with one target text each) is only marginally better than multiple-reference (less source text with multiple target texts each). Future research could test these findings in other MT domains.
@@ 9 Potential Risks
Appendix A Medium dataset statistics
@@
Portuguese 4 1,021 3,261 0.31
Swedish 3 1,249 2,591 0.48
Italian 2 1,325 1,912 0.69
Czech 4 1,759 2,886 0.61
Japanese 9 1,779 2,483 0.72
Chinese 7 2,199 3,671 0.60
Norwegian 2 3,011 4,756 0.63
German 16 6,136 8,905 0.69
Russian 27 20,003 34,030 0.59
French 32 32,353 51,052 0.63
Table 5: Composition of the Medium (#source $=50k$) dataset. "#src" is the total number of unique source paragraphs selected, while column "total" is the all unfiltered number of paragraphs. "%used" is "#src" divided by "total".
@@ Appendix B Example of an untranslatable word
Appendix C Table 6: source texts and target texts from Par3 dataset
source language source text human reference
@@ Appendix C Table 6: source texts and target texts from Par3 dataset
Table 6: Examples of source texts and human references from the Medium dataset where there are meaningful variations between human references for each source text.
Appendix D Figure 3: performance difference between different languages
Figure 3: Performance difference between different languages under different MT metrics when fine-tuning with mT5-large (1st row), LLaMA-2-7B (2nd row), and Opus-MT (3rd row). Darker dots are from Medium+High (#source $=68k$) dataset, and lighter dots are from Unfiltered dataset. The bar chart shows the number of training instances in Unfiltered dataset for each language.